OpenStudy (anonymous):

Use L'Hopital's Rule lim (x / x - 1 - 1 / ln x) x-> 0

OpenStudy (anonymous):

\[\lim \frac{ x }{ x-1 }-\frac{ 1 }{ lnx }\] is it yours?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

x-> 0

OpenStudy (anonymous):

why do we have to use l'hopital rule?

OpenStudy (anonymous):

the teacher makes me use it

OpenStudy (anonymous):

you don't have the form which allows you use that rule, to use that rule, you must have the form 0/0, infinitive/infinitive...and so on. but yours is not. in the first part, when you replace x =0 you have 0/-1 =0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Sorry friend, you may wait for another help. I am helpless here, let me ask other for you @terenzreignz ?

OpenStudy (anonymous):

oh ok ths

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

Terenze: help the asker, please. I am hang on

terenzreignz (terenzreignz):

It doesn't look like L'Hopital applies...

terenzreignz (terenzreignz):

limit goes to zero, as @Hoa said... unless your question wasn't written properly. could you post a picture instead? Or draw it using the draw option @Nageer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):